# DyStream: Streaming Dyadic Talking Heads Generation via Flow Matching-based Autoregressive Model

### v1

@@ 1 Introduction
Recent breakthroughs in full-duplex speech generation systems [wang2024full, defossez2024moshi] have paved the way for highly interactive, streaming AI agents. However, these speech-only agents lack a visual presence, limiting their application in areas like education, sales, and virtual companionship. Bridging this gap requires a model that can generate real-time talking-head video from a single image and the streaming dual-track audio source.
@@
Current state-of-the-art approaches [xu2024vasa, zhu2025infp] have primarily focused on exploring offline solutions. These systems, typically based on two-stage, chunk-by-chunk models, achieve good generation quality. However, their design requires the entire audio chunk’s content before generation begin, introduces a fundamental latency problem. This accumulated latency, often hundreds of milliseconds, is acceptable when the agent is speaking but becomes a critical bottleneck when it needs to act as an active listener. A listener must react instantaneously to the user’s speech with subtle non-verbal cues like nods and facial expressions to create a sense of engagement. The inherent delay in chunk-based systems limits their performance on immediate, naturalistic responses.
@@ 2 Related Work
A prominent approach involves finetuning a pre-trained image-to-video (I2V) model to align the generated video content with an audio condition, mainly for achieving accurate lip synchronization [tian2024emo, xu2024hallo, jiang2024loopy, lin2025omnihuman, chen2025hunyuanvideo_avatar, wei2025mocha, zhou2025paha, bigata2025keysync, li2024ditto, hong2025ACTalker, zhong2025FADA, guan2025audcast, yi2025magicinfinite, yang2025megactor, liu2024tango, liu2025video, qin2025versatile, bigata2025keyface, mu2025flap, sun2024uniavatar, wang2025emotivetalk, chen2025echomimic]. Early works in this domain were often based on U-Net architectures from Stable Diffusion [rombach2022ldm]. For instance, EMO [tian2024emo] integrates audio conditions via an additional cross-attention layer, while Loopy [jiang2024loopy] designs a past-frame compression scheme to better support longer video generation. More recent works have shifted towards Transformer-based backbones (DiT), such as Hallo [xu2024hallo], HunyuanVideo-Avatar [chen2025hunyuanvideo_avatar], and MOCHA [wei2025mocha], which largely follow the cross-attention mechanism to inject audio features. While all these methods demonstrate high-quality results with few artifacts and strong generalization, they suffer from high computational costs and slow inference speeds.
@@
Two-stage methods aim to first predict disentangled appearance and motion representations from audio, and then use these representations to warp a source image [zakharov2022megaportraits, drobyshev2024emoportraits, xu2024vasa, zhu2025infp, lu2021lsp, zhou2020makeittalk, zhang2023sadtalker, wang2025omnitalker, qi2025chatanyone, huang2025se4lip, chen2025audio, gong2025MGGTalk, fu2025DAMC, shen2025Audio-Plane, liu2025disenttalk, zhen2025teller, liu2025syncanimation, guo2025arig]. Representative works such as MegaPortraits [zakharov2022megaportraits], VASA [xu2024vasa], and INFP [zhu2025infp] have explored various designs for a robust, identity-agnostic facial motion space. Recent variants further improve flexibility and style controllability, including OmniTalker [wang2025omnitalker], and ChatAnyone [qi2025chatanyone]. For the audio-to-motion stage, they typically adopt a chunk by chunk diffusion strategy. However, the requirement of the entire audio chunk’s content leads to a latency problem. Different with them, we introduce a flow-matching based autoregressive model to guarantee one token latency in theory.
@@
The early approaches in dyadic talking heads generation typically treated conversations as multi-turn exchanges, where the agent acted as either a speaker or listener in each turn. MultiDialog [GoRealTalk] generated speaker behaviors conditioned on the user’s previous utterance, but failed to produce synchronized listener reactions. To jointly model both roles, ViCo-X [zhou2025interactive] proposes a Role Switcher bridging separate generators for speaker and listener, but the explicit switching often caused unnatural transitions. DIM [dim] pretrained a unified model to capture dyadic context but still required manual role assignment during adaptation. AgentAvatar [wang2023agentavatar] synthesized photo-realistic avatars driven by textual descriptions, but its motions lacked precise audio alignment. Recent person-specific studies [sun2024beyond, ng2024audio, huang2024interact, yan2023dialoguenerf] achieve high realism but perform limited on unseen identities. In contrast to these approaches, which often rely on explicit role-switching rules, multiple expert models, or offline processing of dual audio tracks, our method is designed for online, streaming applications. We utilize a single, unified model to generate both speaker and listener behaviors concurrently from dual-track audio inputs.
@@ 3 Methodology
Given a single image and a streaming dyadic audio signal, our method generates streaming talking-head videos. It is based on a two-stage framework. The first stage disentangles appearance and motion from the source image using an off-the-shelf model based on LIA [wang2024lia], which we briefly review in Section 3.1. The second stage implements audio-to-motion generation using a flow matching-based autoregressive model, as detailed in Section 3.2. Furthermore, to facilitate real-time responsiveness, we design a causal audio encoder with a lookahead module (Section 3.3) to minimize its reliance on future audio frames.
@@ 3.1 Motion-aware Autoencoder
The first stage contains an Autoencoder (AE) to disentangle a source image $\mathbf{I}_{s}$ into two distinct representations: a static appearance feature $\mathbf{v}_{app}$ that encodes the subject’s identity, and a dynamic motion feature $\mathbf{m}$ that describes the pose and expression. Conceptually, this AE includes: an appearance encoder ($\mathcal{E}_{app}$), a motion encoder ($\mathcal{E}_{m}$), a flow estimator ($\mathcal{F}$), and a final decoder ($\mathcal{D}_{vae}$). These components are trained jointly in a self-supervised manner, using a video reconstruction task that reconstruct a driving video by combining the appearance from a source image with the motion from the driving video itself. In our implementation, we adopt the architecture and training objective from LivePortrait [guo2024liveportrait] and LIA [wang2024lia].
@@ 3.2 Flow matching-based Autoregressive Model
As shown in Figure 2, the core architecture of our network is a combination of a flow matching head and an autoregressive (AR) model. Motivated by [li2024autoregressive], the AR model regresses the continuous, deterministic feature, and a single frame-level flow matching head is for stochastic modeling. Specifically, the feature from the AR model, along with the denoising timestep and noise, are fed as conditions for the flow matching head. The denoised clean motion latent is then fed to the AR model for next-frame generation.
@@
For the AR model, we adapt the 1D Transformer block design from Wan2.1 [wan2025wan21]. A key change is our use of block-independent projections for timestamp embeddings instead of the original shared projections. The flow matching head is composed of stacked MLPs. For audio conditioning, the dyadic audio feature is from the concatenation of encoded speaker and listener audio features. We adopt a customized Wav2Vec2 [baevski2020wav2vec] as an audio encoder; the details are in Section 3.3. As shown in Figure 2, we leverage linear interpolation to align with the temporal dimension of the motion features, and fuse the audio feature via element-wise addition. We found that this is similar to a cross-attention with a diagonal attention mask, which shows better performance in our case.
@@ 3.3 Make It Realtime with Customized Wav2Vec2
While current models could generate video in a causal frame-by-frame manner in theory, their lipsync performance is limited with a straightforward implementation. A primary reason is that standard audio encoders, such as Wav2Vec2 [baevski2020wav2vec], typically require non-causal context to extract current features. This leads to severe feature degradation when causality is enforced during inference. As shown in Table 3, the model’s audio-lip synchronization improves as the current audio frame can access more future audio frames. To solve this, we customize a causal Wav2Vec2 with a minimal lookahead module, restricting its receptive field to a maximum of $n$ future audio frames.
To implement this causal Wav2Vec2, we introduce two key modifications to eliminate its inherent non-causal operations. First, the standard architecture utilizes GroupNorm in its initial convolutional layer. Since GroupNorm computes statistics across the entire input sequence, it introduces a dependency on future frames. We replace it with LayerNorm, which operates independently at each timestep, thus ensuring causality. Second, a more significant source of non-causality lies in its positional encoding, implemented as a convolution with a large kernel size of $128$. This layer inherently accesses up to $64$ future frames (a 1.28-second lookahead). We therefore remove this convolutional positional encoding entirely and instead integrate Rotary Positional Embeddings (RoPE) [su2024rope] into the Transformer attention layers.
@@
We then decide the number of lookahead frames based on the demands of real-time processing. For instance, systems like the OpenAI Realtime API [openai2024gpt4ocard] deliver audio in 100ms chunks. Our target video generation rate is 25 fps, which corresponds to one video frame per 40ms of audio. By configuring our modified audio encoder to access an additional 60ms or less of future audio, we can generate a video frame immediately upon receiving the first 100ms audio segment, thereby incurring no additional latency. This lookahead is implemented by adjusting the attention mask inside our customized Wav2Vec2.
In contrast, for the listener part, we adopt a different approach. The average human reaction time in a conversation is approximately 300 milliseconds [conversationlatent]. Unlike a speaker, a listener may not require future audio context to determine precise lip movements. Therefore, for the listener’s audio encoder, we implement a purely causal attention mask. This ensures that the model’s responses are generated based only on past and current audio information.
@@ 4.1 Implementation Details
We train the AE using the AdamW optimizer [loshchilov2017Adamw] with a learning rate of 1e-5 and a batch size of 48. The dimension of the latent motion codes is set to 512. The AE is trained for a total of 150,000 iterations on H200.
@@
We apply a dropout of 0.5 to the speaker and listener audio, and 0.1 to the reference anchor. For inference, we employ a 5-step Euler ODE sampler and utilize Classifier-Free Guidance (CFG) [ho2022cfg], as described in Equation 4, with guidance weights set to $w_{all}=1.0$, $w_{s}=0.5$, $w_{l}=0.5$, and $w_{r}=0.5$. The model is trained for 120,000 iterations using the AdamW optimizer with a learning rate of 2e-5 and a global batch size of 64, and costs 12 hours on a single H200.
@@ 4.2 Dataset
We use two types of data for our experiments: public datasets and an additional colloected dataset. Our data preprocessing pipeline is similar to that of LatentSync [li2024latentsync]. The core steps include detecting and cropping face regions from raw videos, followed by filtering out clips with low SyncNet [Chung16a_SyncNet] scores to ensure data quality. All data are finally processed into talking-head-only videos at a resolution of $512\times 512$, matching the input size of our model. We use two public datasets, HDTF [zhang2021hdtf_dataset] and Realtalk [geng2023theRealTalkData]. After processing, our open-source training set contains approximately 50k video clips, each with a duration of about 8 seconds. Our colloected dataset contains approximately 100k clips of human faces, processed using the same pipeline. This dataset is used exclusively as an additional training source to improve model robustness and is not used for evaluation. The AE is trained on both datasets and the Audio-to-Motion Transformer is trained on Open-Source Datasets only. All videos are resampled to 25FPS.
@@ 4.3 Evaluation of Speaker
We focus on automated, no-reference metrics for evaluation in this subsection. Metrics that require a ground-truth video are discussed in our ablation studies. Following the evaluation protocol of recent video generation benchmarks [huang2024vbench], we evaluate our method across four key dimensions: lip-sync quality [Chung16a_SyncNet], content similarity [radford2021learning], image quality [huang2024vbench], motion dynamics [huang2024vbench]. The specific metrics used to quantify each dimension are detailed in Appendix.
We compare DyStream with several representative or state-of-the-art talking-head generation models: SadTalker [zhang2023sadtalker], Real3DPortrait [ye2024real3d], Hallo3 [cui2024hallo3], Sonic [ji2025sonic], and our reproduced INFP [zhu2025infp]. For evaluation, we construct a unified test set by randomly sampling $100$ clips each from the HDTF and our internal datasets, these clips were filtered out from training. All input faces are cropped and resized to $512\times 512$.
@@
Hallo3 [cui2024hallo3] 6.814 0.915 0.638 0.870
Sonic [ji2025sonic] 8.495 0.935 0.626 0.600
SadTalker [zhang2023sadtalker] 6.704 0.965 0.697 0.080
Real3DPortrait [ye2024real3d] 6.811 0.943 0.637 0.030
INFP* [zhu2025infp] 6.894 0.909 0.643 0.950
@@ 4.4 Evaluation of Listener
To quantitatively evaluate the listener performance, we conduct evaluations on the RealTalk test set, which comprises $38$ videos ranging from $3$ to $14$ seconds in length. We compare our method against our own reproduced of INFP. The evaluation relies on facial expression coefficients (Exp) and head poses (Pose) extracted using MediaPipe [lugaresi2019mediapipe]. We employ four key metrics: Fréchet Distance (FD) and Mean Squared Error (MSE) to measure the realism and accuracy of the generated motions, and two metrics to evaluate diversity: Variance (Var) and SI Diversity (SID). For SID, we follow the methodology of [ng2022learning2listen], which assesses the diversity of facial and head motions by applying k-means clustering (with $k=40$ for expression and k=20 for rotation) on ground truth and then calculating the entropy of the resulting histogram of cluster IDs within each sequence.
@@ 4.5 Ablation Study
Hubert [hsu2021hubert] $\infty$ $\infty$ 8.007
Wavlm [chen2022wavlm] $\infty$ $\infty$ 7.719
Wav2vec [schneider2019wav2vec] 0 0 1.761
Mimi [defossez2024moshi] 0 0 5.628
Encodec(24KHz) [defossez2022encodec] 0 0 2.034
@@ 4.6 Wav Encoder for Online Generation
We then discuss the online version, the selection of the audio encoder will influence the performance on downstream tasks. We conducted a comprehensive evaluation of several prominent audio encoders to determine their suitability for our task. The candidates included non-causal encoders: Wav2vec2 [baevski2020wav2vec], Hubert [hsu2021hubert], WavLM [chen2022wavlm], and causal encoders: Wav2vec [schneider2019wav2vec], Mimi [defossez2024moshi], Encodec [defossez2022encodec], and our Rope-Wav2vec2.
@@
    Apply pretrained causal encoders directly. Instread of using Wav2Vec2 [baevski2020wav2vec], we test wav encoder origianlly designed with causal architectures, such as Mimi [defossez2024moshi], Encodec(24KHz) [defossez2022encodec] and Wav2vec [schneider2019wav2vec].

### v2

@@ 1 Introduction
Recent breakthroughs in full-duplex speech generation systems [61, 11] have paved the way for highly interactive, streaming AI agents. However, these speech-only agents lack a visual presence, limiting their application in areas like education, sales, and virtual companionship. Bridging this gap requires a model that can generate real-time talking-head video from a single image and the streaming dual-track audio source.
@@
Current state-of-the-art approaches [78] have primarily focused on exploring offline solutions. These systems, typically based on two-stage, chunk-by-chunk models, achieve good generation quality. However, their design requires the entire audio chunk’s content before generation begin, introduces a fundamental latency problem. This accumulated latency, often hundreds of milliseconds, is acceptable when the agent is speaking but becomes a critical bottleneck when it needs to act as an active listener. A listener must react instantaneously to the user’s speech with subtle non-verbal cues like nods and facial expressions to create a sense of engagement. The inherent delay in chunk-based systems limits their performance on immediate, naturalistic responses.
@@ 2 Related Work
A prominent approach involves finetuning a pre-trained image-to-video (I2V) model to align the generated video content with an audio condition, mainly for achieving accurate lip synchronization [55, 65, 28, 32, 6, 64, 77, 2, 31, 21, 74, 17, 70, 68, 34, 33, 46, 3, 41, 53, 60, 7]. Early works in this domain were often based on U-Net architectures from Stable Diffusion [48]. For instance, EMO [55] integrates audio conditions via an additional cross-attention layer, while Loopy [28] designs a past-frame compression scheme to better support longer video generation. More recent works have shifted towards Transformer-based backbones (DiT), such as Hallo [65], HunyuanVideo-Avatar [6], and MOCHA [64], which largely follow the cross-attention mechanism to inject audio features. While all these methods demonstrate high-quality results with few artifacts and strong generalization, they suffer from high computational costs and slow inference speeds.
@@
Two-stage methods aim to first predict disentangled appearance and motion representations from audio, and then use these representations to warp a source image [13, 12, 66, 78, 38, 76, 71, 63, 45, 23, 4, 16, 14, 50, 35, 73, 36, 19]. Representative works such as MegaPortraits [13], VASA [66], and INFP [78] have explored various designs for a robust, identity-agnostic facial motion space. Recent variants further improve flexibility and style controllability, including OmniTalker [63], and ChatAnyone [45]. For the audio-to-motion stage, they typically adopt a chunk by chunk diffusion strategy. However, the requirement of the entire audio chunk’s content leads to a latency problem. Different with them, we introduce a flow-matching based autoregressive model to guarantee one token latency in theory.
@@
The early approaches in dyadic talking heads generation typically treated conversations as multi-turn exchanges, where the agent acted as either a speaker or listener in each turn. MultiDialog [44] generated speaker behaviors conditioned on the user’s previous utterance, but failed to produce synchronized listener reactions. To jointly model both roles, ViCo-X [75] proposes a Role Switcher bridging separate generators for speaker and listener, but the explicit switching often caused unnatural transitions. DIM [56] pretrained a unified model to capture dyadic context but still required manual role assignment during adaptation. AgentAvatar [58] synthesized photo-realistic avatars driven by textual descriptions, but its motions lacked precise audio alignment. Recent person-specific studies [52, 43, 24, 67] achieve high realism but perform limited on unseen identities. In contrast to these approaches, which often rely on explicit role-switching rules, multiple expert models, or offline processing of dual audio tracks, our method is designed for online, streaming applications. We utilize a single, unified model to generate both speaker and listener behaviors concurrently from dual-track audio inputs.
@@ 3 Methodology
Given a single image and a streaming dyadic audio signal, our method generates streaming talking-head videos. It is based on a two-stage framework. The first stage disentangles appearance and motion from the source image using an off-the-shelf model based on LIA [62], which we briefly review in Section 3.1. The second stage implements audio-to-motion generation using a flow matching-based autoregressive model, as detailed in Section 3.2. Furthermore, to facilitate real-time responsiveness, we design a causal audio encoder with a lookahead module (Section 3.3) to minimize its reliance on future audio frames.
@@ 3.1 Motion-aware Autoencoder
The first stage contains an Autoencoder (AE) to disentangle a source image $\mathbf{I}_{s}$ into two distinct representations: a static appearance feature $\mathbf{v}_{app}$ that encodes the subject’s identity, and a dynamic motion feature $\mathbf{m}$ that describes the pose and expression. Conceptually, this AE includes: an appearance encoder ($\mathcal{E}_{app}$), a motion encoder ($\mathcal{E}_{m}$), a flow estimator ($\mathcal{F}$), and a final decoder ($\mathcal{D}_{vae}$). These components are trained jointly in a self-supervised manner, using a video reconstruction task that reconstruct a driving video by combining the appearance from a source image with the motion from the driving video itself. In our implementation, we adopt the architecture and training objective from LivePortrait [18] and LIA [62].
@@ 3.2 Flow matching-based Autoregressive Model
As shown in Figure 2, the core architecture of our network is a combination of a flow matching head and an autoregressive (AR) model. Motivated by [30], the AR model regresses the continuous, deterministic feature, and a single frame-level flow matching head is for stochastic modeling. Specifically, the feature from the AR model, along with the denoising timestep and noise, are fed as conditions for the flow matching head. The denoised clean motion latent is then fed to the AR model for next-frame generation.
@@
For the AR model, we adapt the 1D Transformer block design from Wan2.1 [57]. A key change is our use of block-independent projections for timestamp embeddings instead of the original shared projections. The flow matching head is composed of stacked MLPs. For audio conditioning, the dyadic audio feature is from the concatenation of encoded speaker and listener audio features. We adopt a customized Wav2Vec2 [1] as an audio encoder; the details are in Section 3.3. As shown in Figure 2, we leverage linear interpolation to align with the temporal dimension of the motion features, and fuse the audio feature via element-wise addition. We found that this is similar to a cross-attention with a diagonal attention mask, which shows better performance in our case.
@@ 3.3 Make It Realtime with Customized Wav2Vec2
While current models could generate video in a causal frame-by-frame manner in theory, their lipsync performance is limited with a straightforward implementation. A primary reason is that standard audio encoders, such as Wav2Vec2 [1], typically require non-causal context to extract current features. This leads to severe feature degradation when causality is enforced during inference. As shown in Table 3, the model’s audio-lip synchronization improves as the current audio frame can access more future audio frames. To solve this, we customize a causal Wav2Vec2 with a minimal lookahead module, restricting its receptive field to a maximum of $n$ future audio frames.
To implement this causal Wav2Vec2, we introduce two key modifications to eliminate its inherent non-causal operations. First, the standard architecture utilizes GroupNorm in its initial convolutional layer. Since GroupNorm computes statistics across the entire input sequence, it introduces a dependency on future frames. We replace it with LayerNorm, which operates independently at each timestep, thus ensuring causality. Second, a more significant source of non-causality lies in its positional encoding, implemented as a convolution with a large kernel size of $128$. This layer inherently accesses up to $64$ future frames (a 1.28-second lookahead). We therefore remove this convolutional positional encoding entirely and instead integrate Rotary Positional Embeddings (RoPE) [51] into the Transformer attention layers.
@@
We then decide the number of lookahead frames based on the demands of real-time processing. For instance, systems like the OpenAI Realtime API [26] deliver audio in 100ms chunks. Our target video generation rate is 25 fps, which corresponds to one video frame per 40ms of audio. By configuring our modified audio encoder to access an additional 60ms or less of future audio, we can generate a video frame immediately upon receiving the first 100ms audio segment, thereby incurring no additional latency. This lookahead is implemented by adjusting the attention mask inside our customized Wav2Vec2.
In contrast, for the listener part, we adopt a different approach. The average human reaction time in a conversation is approximately 300 milliseconds [40]. Unlike a speaker, a listener may not require future audio context to determine precise lip movements. Therefore, for the listener’s audio encoder, we implement a purely causal attention mask. This ensures that the model’s responses are generated based only on past and current audio information.
@@ 4.1 Implementation Details
We train the AE using the AdamW optimizer [37] with a learning rate of 1e-5 and a batch size of 48. The dimension of the latent motion codes is set to 512. The AE is trained for a total of 150,000 iterations on H200.
@@
We apply a dropout of 0.5 to the speaker and listener audio, and 0.1 to the reference anchor. For inference, we employ a 5-step Euler ODE sampler and utilize Classifier-Free Guidance (CFG) [20], as described in Equation 4, with guidance weights set to $w_{all}=1.0$, $w_{s}=0.5$, $w_{l}=0.5$, and $w_{r}=0.5$. The model is trained for 120,000 iterations using the AdamW optimizer with a learning rate of 2e-5 and a global batch size of 64, and costs 12 hours on a single H200.
@@ 4.2 Dataset
We use two types of data for our experiments: public datasets and an additional colloected dataset. Our data preprocessing pipeline is similar to that of LatentSync [29]. The core steps include detecting and cropping face regions from raw videos, followed by filtering out clips with low SyncNet [8] scores to ensure data quality. All data are finally processed into talking-head-only videos at a resolution of $512\times 512$, matching the input size of our model. We use two public datasets, HDTF [72] and Realtalk [15]. After processing, our open-source training set contains approximately 50k video clips, each with a duration of about 8 seconds. Our colloected dataset contains approximately 100k clips of human faces, processed using the same pipeline. This dataset is used exclusively as an additional training source to improve model robustness and is not used for evaluation. The AE is trained on both datasets and the Audio-to-Motion Transformer is trained on Open-Source Datasets only. All videos are resampled to 25FPS.
@@ 4.3 Evaluation of Speaker
We focus on automated, no-reference metrics for evaluation in this subsection. Metrics that require a ground-truth video are discussed in our ablation studies. Following the evaluation protocol of recent video generation benchmarks [25], we evaluate our method across four key dimensions: lip-sync quality [8], content similarity [47], image quality [25], motion dynamics [25]. The specific metrics used to quantify each dimension are detailed in Appendix.
We compare DyStream with several representative or state-of-the-art talking-head generation models: SadTalker [71], Real3DPortrait [69], Hallo3 [9], Sonic [27], and our reproduced INFP [78]. For evaluation, we construct a unified test set by randomly sampling $100$ clips each from the HDTF and our internal datasets, these clips were filtered out from training. All input faces are cropped and resized to $512\times 512$.
@@
Hallo3 [9] 6.814 0.915 0.638 0.870
Sonic [27] 8.495 0.935 0.626 0.600
SadTalker [71] 6.704 0.965 0.697 0.080
Real3DPortrait [69] 6.811 0.943 0.637 0.030
INFP* [78] 6.894 0.909 0.643 0.950
@@ 4.4 Evaluation of Listener
To quantitatively evaluate the listener performance, we conduct evaluations on the RealTalk test set, which comprises $38$ videos ranging from $3$ to $14$ seconds in length. We compare our method against our own reproduced of INFP. The evaluation relies on facial expression coefficients (Exp) and head poses (Pose) extracted using MediaPipe [39]. We employ four key metrics: Fréchet Distance (FD) and Mean Squared Error (MSE) to measure the realism and accuracy of the generated motions, and two metrics to evaluate diversity: Variance (Var) and SI Diversity (SID). For SID, we follow the methodology of [42], which assesses the diversity of facial and head motions by applying k-means clustering (with $k=40$ for expression and k=20 for rotation) on ground truth and then calculating the entropy of the resulting histogram of cluster IDs within each sequence.
@@ 4.5 Ablation Study
Hubert [22] $\infty$ $\infty$ 8.007
Wavlm [5] $\infty$ $\infty$ 7.719
Wav2vec [49] 0 0 1.761
Mimi [11] 0 0 5.628
Encodec(24KHz) [10] 0 0 2.034
@@ 4.6 Wav Encoder for Online Generation
We then discuss the online version, the selection of the audio encoder will influence the performance on downstream tasks. We conducted a comprehensive evaluation of several prominent audio encoders to determine their suitability for our task. The candidates included non-causal encoders: Wav2vec2 [1], Hubert [22], WavLM [5], and causal encoders: Wav2vec [49], Mimi [11], Encodec [10], and our Rope-Wav2vec2.
@@
    Apply pretrained causal encoders directly. Instread of using Wav2Vec2 [1], we test wav encoder origianlly designed with causal architectures, such as Mimi [11], Encodec(24KHz) [10] and Wav2vec [49].